# Don’t Half-listen: Capturing Key-part Information in Continual Instruction Tuning

### v1

@@ Abstract
Don’t Half-listen: Capturing Key-part Information in
Continual Instruction Tuning
@@ 2.1 Instruction Tuning
LLMs show powerful emergent abilities in many downstream tasks Chowdhery et al. (2023); Touvron et al. (2023b); Zhao et al. (2023). Since most LLMs are typically pretrained with the next word prediction error on large corpora, instruction tuning is proposed as an effective technique to further enhance the instruction-following ability of the generation process Ouyang et al. (2022); Zhang et al. (2023b). And increasing the quantity, diversity, and creativity of instructions is empirically validated as an effective strategy to improve the instruction-following and generalization capabilities of LLMs Zhang et al. (2023b); Xu et al. (2023); Zeng et al. (2023a). Collecting existing datasets and synthesizing data with LLMs are main strategies to obtain high-quality instruction data Zhang et al. (2023b, a). The former collects existing data and converts it into instruction-style datasets through templates or machine translation BELLEGroup (2023); Taori et al. (2023), while the latter like Evol-Instruct Xu et al. (2023), instructWild Ni et al. (2023) and Self-Instruct Wang et al. (2023) ask LLMs to rewrite seed instructions based on specific strategies. In addition to increasing the diversity of task data, optimizing LLM with a comparison dataset collected by human feedback or LLMs also helps to generate desired responses Ouyang et al. (2022); Zhang et al. (2023b). However, LLMs may only remember surface-level patterns of seen instructions, causing the output results not satisfy all constraints on held-out instructions Zhang et al. (2023b); Kung and Peng (2023). In this paper, we propose a key-part information mask mechanism to make LLMs focus more on tokens in instructions that are pertinent to the content, length, and format of the ground truths.
@@ 3.3 Information Gain Fine-tuning
where ${\rm min}$ represents the scaling of $\mathcal{G}(d_{\tau},d_{\tau}^{m})$ into the range $(-\infty,1]$. The lower the information gain, the greater the opportunity we give other tokens to improve the generalization ability of the LLM. By doing so, $\mathcal{L}_{CE}$ maximizes likelihood for complete instances, and $\mathcal{L}_{JSD}$ dynamically adjusts the degree of conservatism when instructions are masked, enabling the LLM to be sensitive to key parts and alleviate the half-listening problem. The detailed implementation is shown in Appendix B.
@@ 3.4 Evaluation Protocol
We evaluate our method on SupNatInst and Domain datasets, where each task contains a task definition, a few demonstrations, and several instances. SupNatInst consists of over 1000 NLP tasks and 76 categories (e.g., text classification, information extraction and etc.) Wang et al. (2022). We select 128 tasks in 40 categories from SupNatInst, 88 tasks are used for training $\mathcal{T}_{seen}$ and 40 as held-out tasks $\mathcal{T}_{unseen}$. And our Chinese domain dataset has 20 tasks and 12 categories, where 13 tasks are used for training $\mathcal{T}_{seen}$ and 7 as held-out tasks $\mathcal{T}_{unseen}$. We use two strategies, single-task (ST) and single-category (SC), to build multi-step training datasets. For the ST setting, we fix $n$ equal to 1, where only 1 task in $\mathcal{T}_{seen}^{t}$ at time step $t$. For the SC setting, we divide seen tasks into multiple groups according to their categories, and train different categories at each time step, because real training scenarios may gradually enhance model abilities of specific categories when training tasks are not available synchronously. Furthermore, to enhance the balance and diversity of each test dataset while accelerating the evaluation process, we sample a few instances for each $\tau_{test}$ based on Self-BLEU Zhu et al. (2018) score and label distribution. More details about datasets can be found in Appendix C.
@@
Table 1: Performance of different methods on Sup-NatInst and Domain datasets. $\ast$ indicates the best, and __\__ indicates the second best. The higher the P-score, the better the model performance. The lower the V-score, the stronger the instruction-following ability. Since INIT usually serves as an upper bound for held-out tasks, and MULTI is usually the upper bound for seen tasks, we report their results.
@@ 4.1 Experimental Setup
We compare our method in the CIT setting with the following baselines. INIT is the foundation LLM without training. MULTI shuffles instances in all training tasks and trains them together. SFT Ouyang et al. (2022) directly fine tunes the LLM on seen tasks sequentially. LoRA Hu et al. (2022) updates the low-rank matrices while the LLM backbone is fixed. L2 and EWC Kirkpatrick et al. (2017) mitigate forgetting by regularizing the loss to penalize the changes of important parameters. DARE Yu et al. (2023) and LM-Cocktail Xiao et al. (2023) obtain the target LLM by model merging, which train multiple models on different tasks and merge them into a single model through weighted average. DYNAINST Mok et al. (2023) determines which instances are stored and replayed based on the entropy of model predictions. PCLL Zhao et al. (2022) and DCL Zeng et al. (2023b) generate pseudo samples for history tasks and utilize knowledge distillation strategy to mitigate catastrophic forgetting.
@@
We choose LLaMA-2-7B-Chat Touvron et al. (2023b) and baichuan-vicuna-chinese-7b  [^1] https://huggingface.co/fireballoon/baichuan-vicuna-chinese-7b as foundation models for experiments on SupNatInst and Domain datasets respectively. Our experiments are implemented based on DeepSpeed Rasley et al. (2020) and FastChat Zheng et al. (2023). And $8$ NVIDIA A100 GPUs are used. We optimize the model parameters by using AdamW optimizer (Loshchilov and Hutter, 2018) with the learning rate of $52e-$ 5. The batch size is $384$ with $16$ gradient accumulation steps and $3$ sentences per GPU. We conduct a grid search to find other hyperparameters that maximize the average P-score on seen and held-out tasks. The optimal settings are: $\{\alpha=0.3,\lambda=0.02,M=10,N=10,epoch=1\}$ on Sup-NatInst and $\{\alpha=0.6,\lambda=0.01,M=3,N=100,epoch=1\}$ on Domain. Additionally, we iteratively perform $30$ evolutions for each task in the instruction-diversity stage. And when evaluating held-out tasks, we add $2$ additional pre-written demonstrations to the input context. More detailed information about the implementation can be found in Appendix B.
@@ 4.2 Main Results
Thirdly, data-replay methods perform better on held-out tasks, and model-merge methods perform better on seen tasks. Compared with DYNAINST, DARE and LM-Cocktail perform better on seen tasks because they selectively inherit abilities of different task categories from multiple models, but abilities are limited when faced with held-out tasks.
@@
Finally, our proposed KPIG achieves the best performance, especially on held-out tasks and the instruction-following ability (V-score). The P-score of KPIG on seen tasks is slightly higher than the model-merge methods while V-score is much lower. And for the held-out tasks, our method performs significantly better than other CL baselines in both P-score and V-score, which shows stronger generalization ability and instruction-following ability. Moreover, the V-score of other baselines on the Domain dataset is much larger than the Sup-NatInst dataset, while our method maintains lower V-score on both Sup-NatInst and domain-specific datasets. This indicates half-listening and instruction violations may be more likely to occur on a specific domain, and our method can better capture the task-aware information and improve the instruction-following ability.
@@ 5 Conclusion
In this paper, we propose a novel CIT paradigm to alleviate problems such as catastrophic forgetting and half-listening, which aims to improve instruction-following ability and generalization ability of LLMs. Our experiments are conducted with the publicly available Super-NaturalInstructions dataset, our in-house dataset, and LLMs from open sources, one of whose initial intentions is to promote the development of instruction-based LLMs. Since LLMs trained with web data may produce toxic content, we must state that the texts generated by our method do not represent our opinions. To alleviate such potential negative impacts, we can adopt appropriate detoxification strategies and principle constraints as default key parts in our method, and we encourage future work to explore these issues.
@@ Appendix A Templates
Appendix B Implementation
@@
[…]
As for the training time, our KPIG takes $340$ minutes to complete the training on Sup-NatInst-ST dataset. Compared with SFT ($200$ minutes), the extra time cost is mainly in the calculation stage of information gain, which takes about $1$ minutes for each time step. In addition, under the setting of $M=10$ and $N=10$, replay-based methods like PCLL, DCL and DYNAINST take about $300$ minutes, and our time difference ($40$ minutes) is that the logits of the masked part need further calculation. However, to achieve the results reported in Table 1, PCLL and DCL need to replay all historical tasks, which takes about $400$ minutes. The above analysis shows that our method offers a relatively balanced trade-off between performance and training efficiency under the setting of CIT.
@@ Appendix C Dataset
Appendix D 6 Held-out Instructions
[…]
Appendix E Details in Human Annotation

### v3

@@ Abstract
Don’t Half-listen: Capturing Key-part Information in Continual Instruction Tuning
@@ 2.1 Instruction Tuning
LLMs show powerful emergent abilities in many downstream tasks Chowdhery et al. (2023); Touvron et al. (2023b); Zhao et al. (2023). Since most LLMs are typically pretrained with the next word prediction error on large corpora, instruction tuning is proposed as an effective technique to further enhance the instruction-following ability of the generation process Ouyang et al. (2022); Zhang et al. (2023b). And increasing the quantity, diversity, and creativity of instructions is empirically validated as an effective strategy to improve the instruction-following and generalization capabilities of LLMs Zhang et al. (2023b); Xu et al. (2023); Zeng et al. (2023a). Collecting existing datasets and synthesizing data with LLMs are main strategies to obtain high-quality instruction data Zhang et al. (2023b, a). The former collects existing data and converts it into instruction-style datasets through templates or machine translation BELLEGroup (2023); Taori et al. (2023), while the latter like Evol-Instruct Xu et al. (2023), instructWild Ni et al. (2023) and Self-Instruct Wang et al. (2023b) ask LLMs to rewrite seed instructions based on specific strategies. In addition to increasing the diversity of task data, optimizing LLM with a comparison dataset collected by human feedback or LLMs also helps to generate desired responses Ouyang et al. (2022); Zhang et al. (2023b). However, LLMs may only remember surface-level patterns of seen instructions, causing the output results not satisfy all constraints on held-out instructions Zhang et al. (2023b); Kung and Peng (2023). In this paper, we propose a key-part information mask mechanism to make LLMs focus more on tokens in instructions that are pertinent to the content, length, and format of the ground truths.
@@ 3.3 Information Gain Fine-tuning
where ${\rm min}$ represents the scaling of $\mathcal{G}(d_{\tau},d_{\tau}^{m})$ into the range $(-\infty,1]$. The lower the information gain, the greater the opportunity we give other tokens to improve the generalization ability of the LLM. By doing so, $\mathcal{L}_{CE}$ maximizes likelihood for complete instances, and $\mathcal{L}_{JSD}$ dynamically adjusts the degree of conservatism when instructions are masked, enabling the LLM to be sensitive to key parts and alleviate the half-listening problem. The detailed implementation is shown in Appendix C.
@@ 3.4 Evaluation Protocol
We evaluate our method on SupNatInst and Domain datasets, where each task contains a task definition, a few demonstrations, and several instances. SupNatInst consists of over 1000 NLP tasks and 76 categories (e.g., text classification, information extraction and etc.) Wang et al. (2022). We select 128 tasks in 40 categories from SupNatInst, 88 tasks are used for training $\mathcal{T}_{seen}$ and 40 as held-out tasks $\mathcal{T}_{unseen}$. And our Chinese domain dataset has 20 tasks and 12 categories, where 13 tasks are used for training $\mathcal{T}_{seen}$ and 7 as held-out tasks $\mathcal{T}_{unseen}$. We use two strategies, single-task (ST) and single-category (SC), to build multi-step training datasets. For the ST setting, we fix $n$ equal to 1, where only 1 task in $\mathcal{T}_{seen}^{t}$ at time step $t$. For the SC setting, we divide seen tasks into multiple groups according to their categories, and train different categories at each time step, because real training scenarios may gradually enhance model abilities of specific categories when training tasks are not available synchronously. Furthermore, to enhance the balance and diversity of each test dataset while accelerating the evaluation process, we sample a few instances for each $\tau_{test}$ based on Self-BLEU Zhu et al. (2018) score and label distribution. More details about datasets can be found in Appendix D.
@@
Table 1: Performance of different methods on Sup-NatInst and Domain datasets. $\ast$ indicates the best, and __\__ indicates the second best. The higher the P-score, the better the model performance. The lower the V-score, the stronger the instruction-following ability. Since INIT sometimes serves as an upper bound for held-out tasks, and MULTI is sometimes the upper bound for seen tasks, we report their results.
@@ 4.1 Experimental Setup
We compare our method in the CIT setting with the following baselines. INIT is the foundation LLM without training. MULTI shuffles instances in all training tasks and trains them together. SFT Ouyang et al. (2022) directly fine tunes the LLM on seen tasks sequentially. LoRA Hu et al. (2022) updates the low-rank matrices while the LLM backbone is fixed. L2 and EWC Kirkpatrick et al. (2017) mitigate forgetting by regularizing the loss to penalize the changes of important parameters. DARE Yu et al. (2023) and LM-Cocktail Xiao et al. (2023) obtain the target LLM by model merging, which train multiple models on different tasks and merge them into a single model through weighted average. DYNAINST Mok et al. (2023) and InsCL Wang et al. (2024) dynamically determines which instances are stored and replayed based on their proposed metrics. PCLL Zhao et al. (2022) and DCL Zeng et al. (2023b) generate pseudo samples for history tasks and utilize knowledge distillation strategy to mitigate catastrophic forgetting.
@@
We choose LLaMA-2-7B-Chat Touvron et al. (2023b) and baichuan-vicuna-chinese-7b  [^1] https://huggingface.co/fireballoon/baichuan-vicuna-chinese-7b as foundation models for experiments on SupNatInst and Domain datasets respectively. Our experiments are implemented based on DeepSpeed Rasley et al. (2020) and FastChat Zheng et al. (2023). And $8$ NVIDIA A100 GPUs are used. We optimize the model parameters by using AdamW optimizer (Loshchilov and Hutter, 2018) with the learning rate of $52e-$ 5. The batch size is $384$ with $16$ gradient accumulation steps and $3$ sentences per GPU. We conduct a grid search to find other hyperparameters that maximize the average P-score on seen and held-out tasks. The optimal settings are: $\{\alpha=0.3,\lambda=0.02,M=10,N=10,epoch=1\}$ on Sup-NatInst and $\{\alpha=0.6,\lambda=0.01,M=3,N=100,epoch=1\}$ on Domain. Additionally, we iteratively perform $30$ evolutions for each task in the instruction-diversity stage. And when evaluating held-out tasks, we add $2$ additional pre-written demonstrations to the input context. More detailed information about the implementation can be found in Appendix C.
@@ 4.2 Main Results
Thirdly, model-merge methods perform better on seen tasks than on held-out tasks. This may be because they selectively inherit abilities of different task categories from multiple models, but abilities are limited when faced with held-out tasks.
@@
Finally, our proposed KPIG achieves the best performance, especially on held-out tasks and the instruction-following ability (V-score). The P-score of KPIG on seen tasks is slightly lower than the InsCL, which may be related to the replayed number. On the held-out tasks, our method performs significantly better than other CL baselines in both P-score and V-score, which shows stronger generalization ability and instruction-following ability. Moreover, the V-score of other baselines on the Domain dataset is much larger than the Sup-NatInst dataset, while our method maintains lower V-score on both Sup-NatInst and domain-specific datasets. This indicates half-listening and instruction violations may be more likely to occur on a specific domain, and our method can better capture the task-aware information and improve the instruction-following ability.
@@ 5 Conclusion
Acknowledgments
Thanks to my colleagues at Meituan, Xuancheng Huang, Lingxun Meng, and Kejin Chen, for their valuable suggestions on improving this work. We also appreciate the assistance from Hongbin Zhang, Xinyu Wang, and Chen Shi of Nanjing University of Science and Technology. We thank all anonymous reviewers for their valuable comments and suggestions on this work.
@@ Acknowledgments
In this paper, we propose a novel CIT paradigm to alleviate the half-listening problem, which aims to improve instruction-following ability and generalization ability of LLMs. Our experiments are conducted with the publicly available Super-NaturalInstructions dataset, our in-house dataset, and LLMs from open sources, one of whose initial intentions is to promote the development of instruction-based LLMs. Since LLMs trained with web data may produce toxic content, we must state that the texts generated by our method do not represent our opinions. To alleviate such potential negative impacts, we can adopt appropriate detoxification strategies and principle constraints, and we encourage future work to explore these issues.
@@ Appendix A Templates
Appendix B Hyperparameter Sensitivity
[…]
As shown in Table 11, we add more experiments on hyperparameter and conduct further analysis. When fixing $\alpha=0.3$, increasing $\lambda$ from 0.001 to 0.1 shows a clear trade-off. While the P-score of seen tasks decreases from 54.7 to 47.6, the P-score of held-out tasks increases from 37.1 to 44.0, indicating stronger regularization benefits for generalization. Moreover, with fixed $\lambda=0.02$, increasing $\alpha$ from 0.3 to 1.0 causes the P-score of both seen and held-out tasks to decline, suggesting excessive values impair overall performance. This may be because as the generation progresses, the posterior probability of the subsequent tokens increases, requiring a smaller $\alpha$ penalty. Therefore, to achieve a relative optimal balance, we choose $\lambda=0.02$ and $\alpha=0.3$ based on the average P-score on seen and held-out tasks. However, we must admit that our exploration of $\lambda$ and $\alpha$ has limitations, because these two indicators may be related to the alignment tax or other issues of the CL process of the LLM. Dynamically adjusting these two hyperparams according to our IG may be more rigorous.
Figure 5: The impact of $N$ on model performance.
$\lambda$ ($\alpha=0.3$) Seen P Unseen P $\alpha$ ($\lambda=0.02$) Seen P Unseen P
0.001 54.7 37.1 0.05 51.9 41.7
0.01 52.6 40.8 0.1 51.6 42.1
0.02 52.2 42.5 0.3 52.2 42.5
0.05 49.6 43.3 0.6 51.3 41.4
0.1 47.6 44.0 1.0 48.6 38.5
Table 11: Performance on Sup-NatInst-ST dataset with varying $\lambda$ and $\alpha$.
Appendix C Implementation
@@
For the experimental settings of other baselines, we give priority to the hyperparameters reported in their paper. In particular, to be fair and reproduce the better performance of baselines, $M$ and $N$ of DYNAINST adopt the same settings as KPIG, and we set $M$ to be the number of historical tasks and $N$ to be the same as KPIG in PCLL and DCL. For InsCL, we set the maximum replayed number to $200$ (consistent with their publication).
As for the training time, our KPIG takes $340$ minutes to complete the training on Sup-NatInst-ST dataset. Compared with SFT ($200$ minutes), the extra time cost is mainly in the calculation stage of information gain, which takes about $1$ minutes for each time step. In addition, under the setting of $M=10$ and $N=10$, replay-based methods like PCLL, DCL and DYNAINST take about $300$ minutes, and our time difference ($40$ minutes) is that the logits of the masked part need further calculation. InsCL takes about $440$ minutes due to the setting of a larger replay number. However, to achieve the results reported in Table 1, PCLL and DCL need to replay all historical tasks, which takes about $400$ minutes. The above analysis shows that our method offers a relatively balanced trade-off between performance and training efficiency under the setting of CIT.
@@ Appendix D Dataset
Appendix E 6 Held-out Instructions
The six held-out instructions we used in our investigations on the information gain (§4.4) are listed in Table 12. The smcalflow classification task is a seen task, which requires determining whether the sentence is spoken by a user or an agent. Ins6 has the smallest information gain and the worst model performance in Table 3. This may be because it is not concise enough and has more redundant constraints compared with other instructions, which may indicate that our information gain may be helpful in measuring the clarity of the task definition.
Figure 6: Saliency scores of task 1645 and 1664.
Appendix F Investigations on Saliency Score
We conduct further experiments on the saliency score Wang et al. (2023a) of the ground-truth and key-part tokens. The saliency score is a crucial metric that indicates how much attention the model pays to relevant segments of user input. We randomly select two tasks from Sup-NatInst, which have significant constraints on the classification label space and the scope of generated content respectively. As shown in Figure 6, ground truth tokens (e.g., ’S’ in ’Similar’ or ’D’ in ’Dissimilar’) receive a higher saliency score in the last 4 layers compared to baselines. For the INIT foundation LLM, attention flow towards the ground-truth tokens from the key parts is more significant than other parts across most tasks. The above results show that KPIG is more effective at identifying the key parts of user intent, leading to more accurate and reliable performance. However, compared to the classification-task 1645, such phenomenon is not so significant for generation-task 1664. This may be because the importance of context is not that different in such tasks, and it is difficult for GPT-4 to accurately identify the key parts (the instructions for such tasks have fewer constraints and definitions). And we believe these phenomena merit further exploration such as more carefully designed data construction and staged training.
Appendix G Details in Human Annotation